# Mass-feeding of jet-launching white dwarfs in grazing and common envelope evolution

### v1

@@ Abstract
I propose a scenario that allows white dwarfs (WDs) to launch relatively powerful jets when they enter a common envelope evolution (CEE) or experience a grazing envelope evolution (GEE) with a red giant branch star (RGB) or an asymptotic giant branch (AGB) star. In this, still a speculative scenario, the accretion for a time is mainly onto an accretion disk with a radius of $\simeq 1R_{\odot}$ that increases in mass. The accretion disk launches the powerful two opposite jets by releasing gravitational energy, up to several times super-Eddington, as its mass increases. The jets that the disk launches remove high-entropy gas from the disk’s outskirts and the envelope that the WD inflates because of nuclear burning on its surface. The motivations to allow WDs to launch powerful jets are recent findings, from the morphologies of post-CEE planetary nebulae, that jets play a major role in the CEE and the accumulating evidence that jets power luminous red novae by jets, as their morphologies indicate. I strengthen my call to include jets in the simulation and modeling of the CEE, consider the GEE as a phase preceding the CEE in many (but not all) cases, and include jets as a major ingredient in modeling and stimulating all energetic luminous red novae.
@@ 1 Introduction
Many planetary nebulae (PNe) morphologies possess a pair or more structural features on opposite sides of the center. The pairs can be of clumps (also called ansae; e.g., Balick 1987), lobes, bubbles, and ‘ears’ (for some studies with images of many PNs see, e.g., Balick 1987; Chu et al. 1987; Schwarz et al. 1992; Corradi & Schwarz 1995; Sahai & Trauger 1998; Sahai et al. 2007, 2011; Parker et al. 2016; Parker 2022). Most studies of the formation of these structures consider shaping by pairs of jets (e.g., Morris 1987; Soker 1990; Sahai & Trauger 1998; Akashi & Soker 2018; Estrella-Trujillo et al. 2019; Tafoya et al. 2019; Balick et al. 2020; Rechy-García et al. 2020; Clairmont et al. 2022; Danehkar 2022; Moraga Baez et al. 2023; Derlopa et al. 2024; Miranda et al. 2024; Sahai et al. 2024; for an alternative scenario see, e.g., Baan et al. 2021). A large fraction of these studies, as well as other studies of jet-launching by PN progenitors, consider the companion to the asymptotic giant branch (AGB) star to launch the pairs of jets. The companion, a main-sequence star or a white dwarf (WD), accretes material from the AGB envelope, while outside the envelope, or inside it during common envelope evolution (CEE), including grazing envelope evolution (GEE). The same holds for red giant branch (RGB) stellar progenitors of PNe (for PN descendants of RGB stars, see e.g., Hillwig et al. 2017; Sahai et al. 2017; Jones et al. 2020, 2022, 2023).
@@
More than a hundred post-AGB nebulae, pre-PNe, and PNe have a central binary system with a short orbital period, i.e., less than several weeks, indicating that the system underwent a common envelope evolution (CEE) phase (e.g., Miszalski et al. 2019; Orosz et al. 2017; Jones 2020, 2024). Orbital periods of several months might indicate that the system has experienced a GEE. The relation between the post-CEE system and jet-shaped PNe led me to claim (Soker, 2025) that jets are CEE’s most robust observable ingredient. However, the majority of hydrodynamical (and magneto-hydrodynamical) simulations do not include jets that the companion to the AGB or RGB star launches (e.g., a partial list from the last decade Staff et al. 2016; Kuruwita et al. 2016; Ohlmann et al. 2016; Iaconi et al. 2017; Chamandy et al. 2019; Law-Smith et al. 2020; Glanz & Perets 2021a, b; González-Bolívar et al. 2022, 2024; Lau et al. 2022a, b; Bermúdez-Bustamante et al. 2024; Chamandy et al. 2024; Gagnier & Pejcha 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024, 2025). Several studies do simulate jets that the more compact companion launches in CEE, including massive stars with a neutron star or a black hole companion; most of these simulations cover a short time, no more than a few orbital periods, or omit other processes (e.g., Moreno Méndez et al. 2017; Shiber & Soker 2018; López-Cámara et al. 2019; Shiber et al. 2019; López-Cámara et al. 2020; Hillel et al. 2022, 2023; López-Cámara et al. 2022; Zou et al. 2022; Soker 2022; Schreier et al. 2023, 2025; Gurjar et al. 2024; Shiber & Iaconi 2024). These arguments, including the extensive list of CEE simulations, underscore the field’s current state: despite considerable progress and efforts over the last decade, we are still far from accurately simulating the CEE, particularly the inclusion of jets that the companion launches.
@@
In general, the compact companion that enters the envelope of a giant star can be a main-sequence star, a WD, a neutron star, or a black hole. A neutron star or a black hole can accrete at very high rates of $\dot{M}_{a}\gtrsim 10^{-3}M_{\odot}{~{}\rm yr}^{-1}$ and up to $\approx 1M_{\odot}{~{}\rm s}^{-1}$, because the accreted material can cool by neutrino emission (Houck & Chevalier, 1991; Chevalier, 1993, 2012); a black hole can as well accrete the gravitational energy. Main-sequence stars can also accrete mass in a CEE. Observationally, some main-sequence stars in post-CEE binary systems of PNe have larger radii than a main-sequence star of the same mass, suggesting they have accreted mass during the CEE (e.g., Jones et al. 2015). The large momenta of the nebulae of some PNe (e.g., Bujarrabal et al. 2001; Sahai et al. 2008) suggest that a main sequence companion can launch jets in the CEE (e.g., Blackman & Lucchini 2014). Theoretically, recent studies have demonstrated that if accretion occurs via an accretion disk and the jets launched from the disk remove the outer high-entropy layers of the newly accreted envelope of the mass-accreting star, then the main-sequence star does not expand significantly (Bear & Soker 2025; Scolnic et al. 2025). The main-sequence star accretes at a high rate without much expansion, ensuring that the accreted gas can form an accretion disk around the main-sequence star and that the potential well stays deep. These two effects allow the launching of relatively energetic jets during the CEE, at least the early CEE, and a GEE that might precede the CEE.
@@ 3 Building a non-steady state accretion disk
The scenario I propose to allow WDs to launch jets during a GEE and early CEE involves a super-Eddington mass accretion rate onto an unsteady accretion disk at a radius of $R_{d}\simeq 1R_{\odot}$. It is this accretion disk that launches the jets. The processes of building an accretion disk at a super-Eddington rate and the launching of a significant fraction of the inflowing gas are similar in some respects to the formation of an accretion disk at a tidal destruction event (e.g., Qiao et al. 2025), where the accretion disk is also not in a steady state.
@@
I take the specific angular momentum of the accreted gas as in Dori et al. (2023), which with the present variables and envelope density profile reads
@@
The envelope is optically thick, so the outflow in the jets takes the extra energy. The disk’s power is larger than the nuclear burning power on the WD, which for a WD of $M_{\rm WD}=0.6M_{\odot}$ can reach a value of $L_{\rm nuc}\simeq 1.5\times 10^{4}L_{\odot}$ (Hachisu et al., 1999). Most of the inflowing gas is ejected back. For example, if the inflowing rate is 5 times as large, $\dot{M}_{\rm inflow}=5\dot{M}_{d}$, the terminal velocity of the jets, with a mass outflow rate of $\dot{M}_{r,2j}=4\dot{M}_{d}$, had they expand freely and not inside the envelope, would be $v_{\rm j}\simeq 0.5v_{\rm Kep,d}\simeq 170{~{}\rm km}{~{}\rm s}^{-1}$, where $v_{\rm Kep,d}$ is the Keplerian velocity in the disk. The WD accretes at a much lower rate due to nuclear burning on its surface. For a WD of $M_{\rm WD}=0.6M_{\odot}$ the accretion rate is $<10^{-6}M_{\odot}{~{}\rm yr}^{-1}$ (Hachisu et al., 1999). The WD inflates an envelope, but the disk outflow, i.e., the jets, removes the outer zones of the inflated envelope in the scenario I propose.
At times, the accretion rate onto the disk can be larger than the scaling of equation (11). The proposed scenario suggests that within several months, for large AGB and RGB stars, a WD companion that enters the envelope with an existing accretion disk can accrete at a high rate to increase the mass in the accretion disk around it, even though the WD is unable to accrete at that rate. The process itself is limited until the building of a high-pressure from a massive accretion disk that slows down the accretion rate. The outflow from the disk removes the high-entropy gas, leaving a low-entropy gas that maintains and builds the accretion disk. The building process of the disk of an average radius of $R_{d}=1R_{\odot}$ releases an energy of
@@ 4 Summary
This study examined a scenario that would allow a WD companion to an RGB or AGB star to launch energetic jets when it enters the common envelope and spirals inwards in the outer zones of the envelope. The WD itself is limited to accreting mass at a low rate because the nuclear burning of the accreted hydrogen-rich gas inflates an envelope (e.g., Hachisu et al. 1999). The limit is more than five orders of magnitude below the possible mass accretion rate from the giant envelope as given by equation (8). Instead of accretion onto the WD, I propose that for a time, the accretion is onto the accretion disk that blows most of the inflowing gas in a bipolar wind, the jets. Still, the accretion disk mass grows and releases gravitational energy during that period. For typical parameters, the accreted gas from the envelope forms an accretion disk with a typical radius of $R_{d}\simeq 1R_{\odot}$ (equation 10). Even if accreting at a net rate of $\simeq 10\%$ of the accretion rate according to equation (8), the accretion process onto the accretion disk can liberate $E_{d}\simeq 10^{46}{~{}\rm erg}$ (equation 12) and few times more within several months to a few years; this is several times the nuclear burning power of the WD. The accretion process can last longer if the WD performs the GEE, namely, the jets efficiently remove the outer zones of the giant’s envelope (the envelope of the RGB or AGB star), and the WD spends some time grazing the giant. In the proposed scenario, the WD already has an accretion disk when it enters the giant’s envelope.
@@
The accretion scenario I proposed here has speculative parts, like the assumption that the disk can grow in mass (up to a limit) by a super-Eddington accretion (by a factor of several). Numerical simulations of accretion by a WD in a CEE or GEE must confirm this assumed flow structure. My motivations to assume/speculate this flow structure that allows WDs to launch energetic jets at the onset of the CEE, or during a GEE phase that might precede the CEE phase, are arguments that support the major role of jets in the CEE (e.g., Soker 2025), and the accumulating evidence for the powering of ILOTs, including luminous red novae, by jets. There are common processes of jets powering transient events (ILOTs, including luminous red novae) and the shaping of PNe (e.g., Soker & Kashi 2012; Soker 2018). The morphologies of spatially resolved ILOTs show bipolar morphologies that only jets can explain without difficulties; the list of bipolar luminous red novae includes V4332 Sgr (Kamiński et al., 2018), Nova 1670 (Shara et al. 1985; Kamiński et al. 2020, 2021a; Kaminski 2024), and V838 Mon (Chesneau et al. 2014; Kamiński et al. 2021b; Mobeen et al. 2021, 2024, 2025). Also, jets are much more efficient in converting kinetic energy to radiation in ILOTs, like luminous red novae, than equatorial mass loss or merger without jets (e.g., Soker 2024). If the accretion process I propose works, a WD can perform the GEE and form jet-powered ILOTs. I argue that most energetic ILOTs, including luminous red novae, are powered by jets, including those ILOTs that studies attribute to CEE or the onset of a CEE (e.g., Karambelkar et al. 2025 for a recent event). Namely, I claim that the launching of energetic jets accompanies the onset of the CEE of the GEE.
@@
Following this study, I strengthen my call to include jets in the simulation and modeling of the CEE, consider the GEE as a phase preceding the CEE in many (but not all) cases and sometimes replacing the CEE, and include jets as a major ingredient in modeling and stimulating all bright luminous red novae, major eruptions of luminous blue variables, and similar gravitationally powered transients, i.e., ILOTs.
@@ Acknowledgements
A grant from the Pazy Foundation supported this research.

### v2

@@ Abstract
I propose a scenario that allows white dwarfs (WDs) to launch relatively powerful jets when they enter a common envelope evolution (CEE) or experience a grazing envelope evolution (GEE) with a red giant branch star (RGB) or an asymptotic giant branch (AGB) star. In this, still a speculative scenario, the accretion for a time is mainly onto an accretion disk with a radius of $\simeq 1R_{\odot}$ that increases in mass. The accretion disk launches the powerful two opposite jets by releasing gravitational energy, up to several times super-Eddington, as its mass increases. The jets launched by the disk remove high-entropy gas from its outskirts and the envelope that the WD inflates due to nuclear burning on its surface. The motivations to allow WDs to launch powerful jets are recent findings, from the morphologies of post-CEE planetary nebulae, that jets play a major role in the CEE and the accumulating evidence that jets power luminous red novae by jets, as their morphologies indicate. I strengthen my call to include jets in the simulation and modeling of the CEE, consider the GEE as a phase preceding the CEE in many (but not all) cases, and include jets as a major ingredient in modeling and simulating all energetic luminous red novae.
keywords:
@@ 1 Introduction
Many planetary nebulae (PNe) morphologies possess a pair or more structural features on opposite sides of the center. The pairs can be of clumps (also called ansae; e.g., Balick 1987), lobes, bubbles, and ‘ears’ (for some studies with images of many PNs see, e.g., Balick 1987; Chu et al. 1987; Schwarz et al. 1992; Corradi & Schwarz 1995; Sahai & Trauger 1998; Sahai et al. 2007, 2011; Parker et al. 2016; Parker 2022). Most studies of the formation of these structures consider shaping by pairs of jets (e.g., Morris 1987; Soker 1990; Garcia-Segura 1997; Sahai & Trauger 1998; García-Segura & López 2000; García-Segura et al. 2005, 2021, 2022; Akashi & Soker 2018; Estrella-Trujillo et al. 2019; Tafoya et al. 2019; Balick et al. 2020; Rechy-García et al. 2020; Clairmont et al. 2022; Danehkar 2022; Moraga Baez et al. 2023; Derlopa et al. 2024; Miranda et al. 2024; Sahai et al. 2024; for an alternative scenario see, e.g., Baan et al. 2021). A large fraction of these studies, as well as other studies of jet-launching by PN progenitors, consider the companion to the asymptotic giant branch (AGB) star to launch the pairs of jets. The companion, a main-sequence star or a white dwarf (WD), accretes material from the AGB envelope, while outside the envelope, or inside it during common envelope evolution (CEE), including grazing envelope evolution (GEE). The same holds for red giant branch (RGB) stellar progenitors of PNe (for PN descendants of RGB stars, see e.g., Hillwig et al. 2017; Sahai et al. 2017; Jones et al. 2020, 2022, 2023).
@@
More than a hundred post-AGB nebulae, pre-PNe, and PNe have a central binary system with a short orbital period, i.e., less than several weeks, indicating that the system underwent a CEE phase (e.g., Miszalski et al. 2019; Orosz et al. 2017; Jones 2020, 2024). Orbital periods of several months might indicate that the system has experienced a GEE. The relation between the post-CEE system and jet-shaped PNe led me to claim (Soker, 2025) that jets are CEE’s most robust observable ingredient. However, the majority of hydrodynamical (and magneto-hydrodynamical) simulations do not include jets that the companion to the AGB or RGB star launches (e.g., a partial list from the last decade Staff et al. 2016; Kuruwita et al. 2016; Ohlmann et al. 2016; Iaconi et al. 2017; Chamandy et al. 2019; Law-Smith et al. 2020; Glanz & Perets 2021a, b; González-Bolívar et al. 2022, 2024; Lau et al. 2022a, b; Bermúdez-Bustamante et al. 2024; Chamandy et al. 2024; Gagnier & Pejcha 2024; Landri et al. 2024; Rosselli-Calderon et al. 2024; Vetter et al. 2024, 2025). Several studies do simulate jets that the more compact companion launches in CEE, including massive stars with a neutron star or a black hole companion; most of these simulations cover a short time, no more than a few orbital periods, or omit other processes (e.g., Moreno Méndez et al. 2017; Shiber & Soker 2018; López-Cámara et al. 2019; Shiber et al. 2019; López-Cámara et al. 2020; Hillel et al. 2022, 2023; López-Cámara et al. 2022; Zou et al. 2022; Soker 2022; Schreier et al. 2023, 2025; Gurjar et al. 2024; Shiber & Iaconi 2024). These arguments, including the extensive list of CEE simulations, underscore the field’s current state: despite considerable progress and efforts over the last decade, we are still far from accurately simulating the CEE, particularly the inclusion of jets that the companion launches.
@@
In general, the compact companion that enters the envelope of a giant star can be a main-sequence star, a WD, a neutron star, or a black hole. A neutron star or a black hole can accrete at very high rates of $\dot{M}_{a}\gtrsim 10^{-3}M_{\odot}~\rm{yr}^{-1}$ and up to $\approx 1M_{\odot}~\rm{s}^{-1}$, because the accreted material can cool by neutrino emission (Houck & Chevalier, 1991; Chevalier, 1993, 2012); a black hole can as well accrete the gravitational energy. Main-sequence stars can also accrete mass in a CEE. Observationally, some main-sequence stars in post-CEE binary systems of PNe have larger radii than a main-sequence star of the same mass, suggesting they have accreted mass during the CEE (e.g., Jones et al. 2015). The large momenta of the nebulae of some PNe (e.g., Bujarrabal et al. 2001; Sahai et al. 2008) suggest that a main-sequence companion can launch jets in the CEE (e.g., Blackman & Lucchini 2014). Theoretically, recent studies have demonstrated that if accretion occurs via an accretion disk and the jets launched from the disk remove the outer high-entropy layers of the newly accreted envelope of the mass-accreting star, then the main-sequence star does not expand significantly (Bear & Soker 2025; Scolnic et al. 2025). The main-sequence star accretes at a high rate without much expansion, ensuring that the accreted gas can form an accretion disk around the main-sequence star and that the potential well stays deep. These two effects allow the launching of relatively energetic jets during the CEE, at least the early CEE, and a GEE that might precede the CEE.
@@ 3 Building a non-steady state accretion disk
The scenario I propose to allow WDs to launch jets during a GEE and early CEE involves a super-Eddington mass accretion rate onto an unsteady accretion disk at a radius of $R_{d}\simeq 1R_{\odot}$. It is this accretion disk that launches the jets. The processes of building an accretion disk at a super-Eddington rate and the launching of a significant fraction of the inflowing gas are similar in some respects to the formation of an accretion disk at a tidal destruction event (e.g., Qiao et al. 2025), where the accretion disk is also not in a steady state. Figure 1 presents a cartoon of the proposed scenario.
Figure 1: A schematic presentation of the proposed scenario (not to scale) in the meridional plane; there is an axial symmetry around an axis through the WD and perpendicular to the accretion disk. The main energy source of the jets is the gravitational energy due to the mass increase of the accretion disk. The WD accretes at a rate below the one that leads to envelope inflation.
@@
I take the specific angular momentum of the accreted gas as in Dori et al. (2023), which, with the present variables and envelope density profile, reads
@@
The envelope is optically thick, so the outflow in the jets takes the extra energy. The disk’s power is larger than the nuclear burning power on the WD, which for a WD of $M_{\rm WD}=0.6M_{\odot}$ can reach a value of $L_{\rm nuc}\simeq 1.5\times 10^{4}L_{\odot}$ (Hachisu et al., 1999). Most of the inflowing gas is ejected back. For example, if the inflowing rate is 5 times as large, $\dot{M}_{\rm inflow}=5\dot{M}_{d}$, the terminal velocity of the jets, with a mass outflow rate of $\dot{M}_{r,2j}=4\dot{M}_{d}$, had they expand freely and not inside the envelope, would be $v_{\rm j}\simeq 0.5v_{\rm Kep,d}\simeq 170~\rm{km}~\rm{s}^{-1}$, where $v_{\rm Kep,d}$ is the Keplerian velocity in the disk. The WD accretes at a much lower rate due to nuclear burning on its surface. For a WD of $M_{\rm WD}=0.6M_{\odot}$ the accretion rate is $10^{-6}M_{\odot}~\rm{yr}^{-1}$ (Hachisu et al., 1999). The WD inflates an envelope, but the disk outflow, i.e., the jets, removes the outer zones of the inflated envelope in the scenario I propose.
At times, the accretion rate onto the disk can be larger than the scaling of equation (11). The proposed scenario suggests that within several months, for large AGB and RGB stars, a WD companion that enters the envelope with an existing accretion disk can accrete at a high rate to increase the mass in the accretion disk around it, even though the WD is unable to accrete at that rate. The process itself is limited until a high-pressure state is built from a massive accretion disk, which slows down the accretion rate, or the accretion rate onto the WD itself, $\dot{M}_{\rm WD}$, surpasses the value above which the WD inflates an envelope. In the latter case, the inflated envelope engulfs the accretion disk and prevents further accretion onto it, shutting down the accretion process via a disk. The disk itself might survive until it is depleted, but no further accretion onto a disk takes place.
The outflow from the disk removes the high-entropy gas, leaving a low-entropy gas that maintains and builds the accretion disk. The building process of the disk of an average radius of $R_{d}=1R_{\odot}$ releases an energy of
@@ 4 Summary
This study examined a scenario that would allow a WD companion to an RGB or AGB star to launch energetic jets when it enters the common envelope and spirals inwards in the outer zones of the envelope. The WD itself is limited to accreting mass at a low rate because the nuclear burning of the accreted hydrogen-rich gas inflates an envelope (e.g., Hachisu et al. 1999). The limit is more than five orders of magnitude below the possible mass accretion rate from the giant envelope as given by equation (8). Instead of accretion onto the WD, I propose that for a time, the accretion occurs onto the accretion disk, which then blows most of the inflowing gas in a bipolar wind, forming the jets. Still, the accretion disk mass grows and releases gravitational energy during that period (Figure 1). For typical parameters, the accreted gas from the envelope forms an accretion disk with a typical radius of $R_{d}\simeq 1R_{\odot}$ (equation 10). Even if accreting at a net rate of $\simeq 10\%$ of the accretion rate according to equation (8), the accretion process onto the accretion disk can liberate $E_{d}\simeq 10^{46}~\rm{erg}$ (equation 12) and few times more within several months to a few years; this is several times the nuclear burning power of the WD. The accretion process can last longer if the WD performs the GEE, namely, the jets efficiently remove the outer zones of the giant’s envelope (the envelope of the RGB or AGB star), and the WD spends some time grazing the giant. In the proposed scenario, the WD already has an accretion disk when it enters the giant’s envelope.
@@
The accretion scenario I proposed here has speculative parts, like the assumption that the disk can grow in mass (up to a limit) by a super-Eddington accretion (by a factor of several). Numerical simulations of accretion by a WD in a CEE or GEE must confirm this assumed flow structure. As well, the simulations should include the effect of magnetic fields, as the flow develops a dynamo that amplifies magnetic fields (e.g., Nordhaus et al. 2011). My motivations to assume/speculate this flow structure that allows WDs to launch energetic jets at the onset of the CEE, or during a GEE phase that might precede the CEE phase, are arguments that support the major role of jets in the CEE (e.g., Soker 2025), and the accumulating evidence for the powering of ILOTs, including luminous red novae, by jets. There are common processes of jets powering transient events (ILOTs, including luminous red novae) and the shaping of PNe (e.g., Soker & Kashi 2012; Soker 2018). The morphologies of spatially resolved ILOTs show bipolar morphologies that only jets can explain without difficulties; the list of bipolar luminous red novae includes V4332 Sgr (Kamiński et al., 2018), Nova 1670 (Shara et al. 1985; Kamiński et al. 2020, 2021a; Kaminski 2024), and V838 Mon (Chesneau et al. 2014; Kamiński et al. 2021b; Mobeen et al. 2021, 2024, 2025). Also, jets are much more efficient in converting kinetic energy to radiation in ILOTs, like luminous red novae, than equatorial mass loss or merger without jets (e.g., Soker 2024). If the accretion process I propose works, a WD can perform the GEE and form jet-powered ILOTs. I argue that most energetic ILOTs, including luminous red novae, are powered by jets, including those ILOTs that studies attribute to CEE or the onset of a CEE (e.g., Karambelkar et al. 2025 for a recent event). Namely, I claim that the launching of energetic jets accompanies the onset of the CEE or the GEE.
@@
Following this study, I strengthen my call to include jets in the simulation and modeling of the CEE, consider the GEE as a phase preceding the CEE in many (but not all) cases and sometimes replacing the CEE, and include jets as a major ingredient in modeling and simulating all bright luminous red novae, major eruptions of luminous blue variables, and similar gravitationally powered transients, i.e., ILOTs.
@@ Acknowledgements
I thank an anonymous referee for useful suggestions. I thank the Charles Wolfson Academic Chair at the Technion for the support. A grant from the Pazy Foundation supported this research.